Divide 1290 into A,B and C such that A is 2/5 of B and B:C = 4:3.
Unable to solve please help.

Found 2 solutions by Boreal, jorel1380:
Answer by Boreal(15235)

a is 2/5 th of B
so a=2:b=5
where b=4, c=3
so where b=5, c=3.75 (set it up as a proportion and cross-multiply)
2:5:3.75
multiply these by 4 to make them integers
8:20:15; note that A:B is 2:5 and B:C as 4:3
add these 3 (43)
divide 1290/43=30
that ratio, multiplied by 30, will be the answer
240:600:450
They add to 1290.

Answer by jorel1380(3719)

A+B+C=1290
A=2B/5
B/C=4/3,
C=3B/4
2B/5+B+3B/4=1290
8B+20B+15B=25800
B=600
A=240
C=450
☺☺☺☺